# Conceiving Naturally After IVF:

### v1

@@ Introduction
Indeed, while the characteristics of couples who use ART strongly correlate with poorer outcomes at birth111 This is not only true for multiple pregnancies, but also for singleton pregnancies. Singleton pregnancies are pregnancies where the mother is expecting only one child. Multiple pregnancies, such as twins and triplets, increase risks of adverse outcomes at birth for the mother and the children. They are common among IVF pregnancies although largely preventable by transferring a single embryo per each cycle. The Reproductive Technology Accreditation Committee of the Fertility Society of Australia and New Zealand requires to “minimise the incidence of multiple pregnancies”, which includes recommending a single embryo transfer per cycle, ensuring that no more than two embryos are transferred per cycle, and informing the patient of the risks associated with double transfers (Reproductive Technology Accreditation Committee, 2021). , both for mother (e.g. caesarean sections) and the baby (e.g. pre-term births), it remains unclear whether ART treatment independently increases the risk of such adverse outcomes. These characteristics include, for instance, advanced maternal age (Attali and Yogev, 2021; Lean et al., 2017; Berntsen et al., 2019), certain causes of infertility and comorbidities including obesity (Pasquali et al., 2007; Maheshwari et al., 2007). In relation to obstetric interventions, one concern is that the considerable emotional and financial investment implied by an ART pregnancy might exacerbate defensive medical behaviour, where physicians intervene more than necessary driven by the fear of medical liability (for evidence of its impact on c-section rates, see Dubay et al., 1999; Jena et al., 2015; Betrán et al., 2018; Bertoli and Grembi, 2019; Card et al., 2019; Laurita Longo et al., 2020).
@@
In this paper, we study the effect of ART treatment on obstetric and perinatal outcomes. We measure its impact on the probability of pre-term birth and, more broadly, on the health of the baby. This includes the impact of ART on the baby’s gestational age at birth, weight at birth and its APGAR 1 and 5 scores (a measure of infant health). Finally, we study its impact on the rates of spontaneous labour (rather than induced), preterm spontaneous labour and c-section, which helps us understand whether ART pregnancies experience higher rates of interventions, holding all else constant.
@@
Our approach leverages the stochastic nature of conception, including ART conception. Indeed, while factors such as age and reproductive health influence the probability of conception, chance (idiosyncratic factors) plays an important role. Fertile couples in their 20s have a $25\%$ chance of conceiving spontaneously in each menstrual cycle (ACOG, 2023), and a $41-47\%$ probability of conceiving via IVF per IVF cycle222 Estimates computed on 7 February 2024 via the online government-funded IVF prediction tools: YourIVFSuccess.com.au, based on Australian data, https://www.cdc.gov/art/ivf-success-estimator/index.html, based on US data, and https://w3.abdn.ac.uk/clsm/opis/tool/ivf1, based on UK data. The parameters set for the first prediction tool were: “Age of person intending to carry the pregnancy: 22”, “Sperm provider’s age: 22”, “Previous pregnancy: No”, “Main infertility diagnosis: I have not had any tests carried out, or I do not have medical infertility”, “Previous IVF treatment: No”. The outcome was ”having a baby in the 1st complete egg retrieval cycle” For the second tool the parameters were: “Age: 22 years”, “Weight: 110 lbs”, “Height: 5 feet”, “Number of IVF cycles used: I’ve never used IVF”, “Number of prior pregnancies: None ”, “Number of prior births: None”, “Egg source: My own eggs”, “Diagnosis: I don’t know/no reason”. The outcome was “cumulative chance of live birth after 1 retrieval and all transfers within 12 months.” For the third tool, the parameters were: “What is your age? 22”, “How many years have you been trying to conceive? 0”, “Have you been pregnant before? No”, “Do you have a problem with your tubes? No”, “Do you have an ovulation problem? No”, “Do you have a male factor fertility problem? No”, “Do you have an unexplained fertility problem? No”, “Do you plan to have IVF or ICSI? IVF” The outcome was “chance of having your first baby after 1 complete cycle of treatment.” . In either case, the residual variation in conception remains large and is plausibly idiosyncratic. The persistently low precision of clinical prediction models for IVF pregnancies, even the most recent ones (C-indices between 0.6-0.7 McLernon and Bhattacharya, 2023), is also consistent with some variation in IVF success being idiosyncratic.
@@
Second, we flexibly adjust for key confounding variables, using both standard linear methods and flexible machine-learning-based methods (see Figure 2). Our linked administrative datasets allow us to observe both cycle-level information and some clinical history of the mother, which we use to control for key confounders such as age, parity, number of prior ART cycles, socio-economic characteristics and co-morbidities. Finally, we perform a number of sensitivity analyses, studying how unobserved confounders might affect our estimates (via Cinelli and Hazlett, 2019) and exploring how our results change using alternative sample restrictions and model specifications.
We do not find evidence that ART increases the risk of preterm birth or other adverse obstetric outcomes. In fact, the effects of ART on gestational age at birth, birth weight and APGAR 1 and 5 scores are not statistically different from zero at conventional levels. Our estimated null effects are precise, so that it is unlikely that clinically significant effect sizes exist and have not been detected 333 We can exclude the following (clinically small) effect sizes for the effect of ART on obstetric outcomes: effects between [-0.16, 0.10] weeks for gestational age at birth; between [-1.1, 1.9] p.p. for the risk of preterm birth; between [-0.07, 0.09] points for the APGAR1 score; between [-0.02, 0.1] points for the APGAR5 score; between [-14, 60] grams for the effect of ART on birth weight. . Moreover, we estimate that ART has a (precise) null effect on the risk of preterm spontaneous birth, an important adverse outcome that is not subject to medical intervention and indicates poor infant health. We also find that ART slightly decreases the risk of obstetric interventions, lowering the risk of a caesarean section (-5.1 p.p.) and increasing the rate of spontaneous labour (+4 p.p.).
@@
Then, we drop births with at least one missing observation, 7477 treatment births (38.1%) and 521 control births (35.8%) leaving us with 12,169 treatment births and 935 control births555 These values change slightly across our outcomes at the estimation stage as, when we run the analysis for each outcome, we drop the births for which that outcome is missing.. These missing rates are mostly driven by the gestational diabetes and diabetes mellitus variables, of which 32% of the observations are missing. These are important variables, either confounders or closing a back-door path, and excluding them would likely bias our estimates. Our other covariates have less than 0.5% of missing observations with the exception of “smoking”, which has 3.6%.
The key issue informing the width of the conception window is that prospective mothers experiencing an unsuccessful ART cycle tend to schedule the next one at a close date, if they decide to schedule one666 In our sample of subfertile and ART-conceived singletons, the median time between unsuccessful cycles is 97 days. The 25th percentile is 56 days and the 75th percentile is 322 days. . Hence, if we compare a mother that has a successful ART cycle with one that conceived spontaneously a long time after a failed cycle, we are comparing a woman that might have undertaken several more cycles, had she been unsuccessful, with a woman that did not do that. These two women are likely different in important ways, and this would likely be a source of selection bias in our analysis. Figure 3 can help visualising how the control group is constructed.
@@
Our outcomes and covariates of interest are presented in Table LABEL:tab:des. These covariates have been selected as they are key confounders777 While pre-eclampsia, gestational hypertension and gestational diabetes are not confounders, it is important to include them as they might be blocking a backdoor path or mediate the effect of ART on obstetric outcomes, the latter meaning that they can both have an impact on our outcomes and on the probability of a successful ART cycle—the treatment.
@@
Second, the confounder recording the number of previous failed ART cycles was based on the number of autologous ART cycles888 An autologous ART cycle is an ART cycle where the woman uses her oocyte or embryo, rather than one donated by a third party. In other words, we excluded donor cycles. , fresh or thaw—excluding freeze-all oocyte and/or embryo cycles 999 These are cycles where a woman freezes her oocytes or embryos for future use. . Women donating or freezing their oocytes and/or eggs were excluded because they weren’t at risk of pregnancy.
@@ Methods
If this this two-step randomisation model were the correct model of the world, we should observe excellent balance between our groups of interest—Group A (spontaneous conception) and Group D (ART conception). We do not impose this strong assumption though. Instead, we make a conditional exogeneity/independence claim. This means assuming that the treatment assignments are independent of our outcomes of interest after controlling for the relevant confounders—variables causing changes in both treatment and outcome.
@@
We start from a descriptive analysis of our sample. Table LABEL:tab:tab:cov_balance_anzard_12 reports difference in means across treatment groups. We find a clinically important difference in maternal age, which is higher by 1.9 years (36.6-34.7 years) in mothers that spontaneously conceived after a failed ART —at an age where maternal fertility is in significant decline. This is also reflected in a lower maternal year of birth. In our sample, mothers who spontaneously conceived after ART are less likely to be at their first cycle ($0.25-0.36=-0.10$ p.p.) and have had more unsuccessful cycles before their pregnancy ($3.1-1.6=1.5$ cycles). A last noticeable difference is in the prevalence of male infertility, which is 7 p.p. higher in mothers of ART-conceived babies. The two groups are clinically similar in other respects, such as number of past pregnancies, socio-economic disadvantage, co-morbidities and cause of infertility.
@@ Results
Table LABEL:tab:tab:ols reports the OLS estimates of Equation 1 under our full sample (column 1), and under further sample restrictions (columns 2-6)  [^10] When restricting the sample, we adjust the model accordingly, so that, for instance, we drop the Number of previous pregnancies variable if we are conditioning on having no previous pregnancies.. In column 2, we only include births whose mother had no previous pregnancies longer than 20 weeks of gestation (nulliparous) nor previous unsuccessful ART cycles. In column 3, we only include births whose mother had no previous pregnancies longer than 20 weeks of gestation. In column 4, we only include, in the control group, births whose mother had a single previous unsuccessful ART cycles, and in the treatment group, ART births conceived at the first ART cycle. Finally, in column 5 we drop all our covariates and instead stratify by our most important covariates. More precisely, we estimate a difference in means across treatment groups in a sample restricted to nulliparous mothers at their first ART cycle, aged less than 40 and without comorbidities (selected primips).
@@
The estimated treatment effects are, in general, clinically small [TO BE CONFIRMED BY THE CLINICIAN]. Focusing on our baseline estimates (Column 1), we find that ART has no significant effect on the rate of preterm births, on the rate of spontaneous preterm births, and on the average gestational age at birth—neither at the statistical nor at the clinical level [TO BE CONFIRMED BY THE CLINICIAN]. Taking preterm birth as an example, the 0.008 coefficient, with a 0.01 standard error, indicates that our best estimate for the effect of ART treatment of preterm birth is 0.8 p.p., but we cannot exclude that the true estimate is equal to zero at the 95% confidence level. The implied confidence interval is [$-0.0116,0.0276$], so that we cannot exclude effects ranging from -1.16 p.p. to +2.76 p.p [^11] Confidence intervals, while omitted in Table LABEL:tab:tab:ols due to space constraints, are reported for the double machine learning estimates, which are very similar to the OLS estimates and are reported Tables LABEL:tab:tab:dml_forest_ate_anzard_12 and LABEL:tab:tab:dml_forest_ate_anzard_12_trim_15.
@@
The same applies to birth weight and to the APGAR 1 and 5 scores, whose estimated effects are not statistically significant and are clinically small [TO BE CONFIRMED BY THE CLINICIAN]. To give an idea of the magnitudes of these statistically insignificant estimates, the effect of ART on APGAR1 is estimated to be $-0.013$ points, on a 1-to-5 scale, while the effect on gestational age at birth is estimated to be $-0.057$ weeks, which less than 10 hours. As shown by the p-values in the square brackets, these estimates are not statistically significant at the 1% level or below, despite small standard errors. Considering that gestational age is measured in weeks, the APGAR scores are integers between zero and five and that birth weight is measured grams, we can notice how imposing additional restrictions on the sample does not shift the estimates in a clinically significant way for the most part [TO BE CONFIRMED BY THE CLINICIAN]. Understandably, our estimates’ standard errors and statistical significance varies as the sample gets smaller.
@@
Table LABEL:tab:tab:dml_forest_ate_anzard_12 reports the baseline DML treatment effect estimates, while Table LABEL:tab:tab:dml_forest_ate_anzard_12_trim_15 reports the DML estimates after trimming the propensity score at 0.15 and 0.85. Not only the baseline DML estimates are robust to trimming, they are also virtually identical to the OLS estimates (Table LABEL:tab:tab:ols , Column 1), as can be visualised in Figures 7 and 6.

### v2

@@ Introduction
Indeed, while the characteristics of couples who use ART strongly correlate with poorer outcomes at birth111 This is not only true for multiple pregnancies, but also for singleton pregnancies. Singleton pregnancies are pregnancies where the mother is expecting only one child. Multiple pregnancies, such as twins and triplets, increase risks of adverse outcomes at birth for the mother and the children. They are common among IVF pregnancies although largely preventable by transferring a single embryo per each cycle. The Reproductive Technology Accreditation Committee of the Fertility Society of Australia and New Zealand requires to “minimise the incidence of multiple pregnancies”, which includes recommending a single embryo transfer per cycle, ensuring that no more than two embryos are transferred per cycle, and informing the patient of the risks associated with double transfers (Reproductive Technology Accreditation Committee, 2021). , both for mother (e.g. caesarean sections) and the baby (e.g. pre-term births), it remains unclear whether ART treatment independently increases the risk of such adverse outcomes. These characteristics include, for instance, advanced maternal age (Attali2012; Lean et al., 2017; Berntsen et al., 2019), certain causes of infertility and comorbidities including obesity (Pasquali et al., 2007; Maheshwari et al., 2007). In relation to obstetric interventions, one concern is that the considerable emotional and financial investment implied by an ART pregnancy might exacerbate defensive medical behaviour, where physicians intervene more than necessary driven by the fear of medical liability (for evidence of its impact on c-section rates, see Dubay et al., 1999; Jena et al., 2015; Betrán et al., 2018; Bertoli and Grembi, 2019; Card et al., 2019; Laurita Longo et al., 2020).
@@
In this paper, we study the effect of ART treatment222 on obstetric and perinatal outcomes. We measure its impact on the probability of pre-term birth and, more broadly, on the health of the baby. This includes the impact of ART on the baby’s gestational age at birth, weight at birth and its APGAR 1 and 5 scores (a measure of infant health). Finally, we study its impact on the rates of spontaneous labour (rather than induced), preterm spontaneous labour and c-section, which helps us understand whether ART pregnancies experience higher rates of interventions, holding all else constant.
@@
Our approach leverages the stochastic nature of conception, including ART conception. Indeed, while factors such as age and reproductive health influence the probability of conception, chance (idiosyncratic factors) plays an important role. Fertile couples in their 20s have a $25\%$ chance of conceiving spontaneously in each menstrual cycle (ACOG, 2023), and a $41-47\%$ probability of conceiving via IVF per IVF cycle333 Estimates computed on 7 February 2024 via the online government-funded IVF prediction tools: YourIVFSuccess.com.au, based on Australian data, https://www.cdc.gov/art/ivf-success-estimator/index.html, based on US data, and https://w3.abdn.ac.uk/clsm/opis/tool/ivf1, based on UK data. The parameters set for the first prediction tool were: “Age of person intending to carry the pregnancy: 22”, “Sperm provider’s age: 22”, “Previous pregnancy: No”, “Main infertility diagnosis: I have not had any tests carried out, or I do not have medical infertility”, “Previous IVF treatment: No”. The outcome was ”having a baby in the 1st complete egg retrieval cycle” For the second tool the parameters were: “Age: 22 years”, “Weight: 110 lbs”, “Height: 5 feet”, “Number of IVF cycles used: I’ve never used IVF”, “Number of prior pregnancies: None ”, “Number of prior births: None”, “Egg source: My own eggs”, “Diagnosis: I don’t know/no reason”. The outcome was “cumulative chance of live birth after 1 retrieval and all transfers within 12 months.” For the third tool, the parameters were: “What is your age? 22”, “How many years have you been trying to conceive? 0”, “Have you been pregnant before? No”, “Do you have a problem with your tubes? No”, “Do you have an ovulation problem? No”, “Do you have a male factor fertility problem? No”, “Do you have an unexplained fertility problem? No”, “Do you plan to have IVF or ICSI? IVF” The outcome was “chance of having your first baby after 1 complete cycle of treatment.” . In either case, the residual variation in conception remains large and is plausibly idiosyncratic. The persistently low precision of clinical prediction models for IVF pregnancies, even the most recent ones (C-indices between 0.6-0.7 McLernon and Bhattacharya, 2023), is also consistent with some variation in IVF success being idiosyncratic.
@@
Second, we adjust for key confounding variables, using both standard linear methods and a flexible approach—double machine learning (DML) with random forests (see Figure 2). Our linked administrative datasets allow us to observe both cycle-level information and some clinical history of the mother, which we use to control for key confounders such as age, parity, number of prior ART cycles, socio-economic characteristics and co-morbidities. Finally, we perform a number of sensitivity analyses, studying how unobserved confounders might affect our estimates (via Cinelli and Hazlett, 2019) and exploring how our results change using alternative sample restrictions and model specifications.
We do not find evidence that ART increases the risk of preterm birth or other adverse obstetric outcomes. In fact, the effects of ART on gestational age at birth, birth weight and APGAR 1 and 5 scores are not statistically different from zero at conventional levels. Our estimated null effects are precise, so that it is unlikely that clinically significant effect sizes exist and have not been detected 444 We can exclude the following (clinically small) effect sizes for the effect of ART on obstetric outcomes: effects between [-0.16, 0.10] weeks for gestational age at birth; between [-1.1, 1.9] p.p. for the risk of preterm birth; between [-0.07, 0.09] points for the APGAR1 score; between [-0.02, 0.1] points for the APGAR5 score; between [-14, 60] grams for the effect of ART on birth weight. . Moreover, we estimate that ART has a (precise) null effect on the risk of preterm spontaneous birth, an important adverse outcome that is not subject to medical intervention and indicates poor infant health. We also find that ART slightly decreases the risk of obstetric interventions, lowering the risk of a caesarean section (-5.1 p.p.) and increasing the rate of spontaneous labour (+4 p.p.).
@@
Then, we drop births with at least one missing observation, 7477 treatment births (38.1%) and 521 control births (35.8%) leaving us with 12,169 treatment births and 935 control births666 These values change slightly across our outcomes at the estimation stage as, when we run the analysis for each outcome, we drop the births for which that outcome is missing.. These missing rates are mostly driven by the gestational diabetes and diabetes mellitus variables, of which 32% of the observations are missing. These are important variables, either confounders or closing a back-door path, and excluding them would likely bias our estimates. Our other covariates have less than 0.5% of missing observations with the exception of “smoking”, which has 3.6%.
The key issue informing the width of the conception window is that prospective mothers experiencing an unsuccessful ART cycle tend to schedule the next one at a close date, if they decide to schedule one777 In our sample of subfertile and ART-conceived singletons, the median time between unsuccessful cycles is 97 days. The 25th percentile is 56 days and the 75th percentile is 322 days. . Hence, if we compare a mother that has a successful ART cycle with one that conceived spontaneously a long time after a failed cycle, we are comparing a woman that might have undertaken several more cycles, had she been unsuccessful, with a woman that did not do that. These two women are likely different in important ways, and this would likely be a source of selection bias in our analysis. Figure 3 can help visualising how the control group is constructed.
@@
Our outcomes and covariates of interest are presented in Table LABEL:tab:des. These covariates have been selected as they are key confounders888 While pre-eclampsia, gestational hypertension and gestational diabetes are not confounders, it is important to include them as they might be blocking a backdoor path or mediate the effect of ART on obstetric outcomes, the latter meaning that they can both have an impact on our outcomes and on the probability of a successful ART cycle—the treatment.
@@
Second, the confounder recording the number of previous failed ART cycles was based on the number of autologous ART cycles999 An autologous ART cycle is an ART cycle where the woman uses her oocyte or embryo, rather than one donated by a third party. In other words, we excluded donor cycles. , fresh or thaw—excluding freeze-all oocyte and/or embryo cycles 101010 These are cycles where a woman freezes her oocytes or embryos for future use. . Women donating or freezing their oocytes and/or eggs were excluded because they weren’t at risk of pregnancy.
@@ Methods
If is this two-step randomisation model were the correct model of the world, we should observe excellent balance between our groups of interest—Group A (spontaneous conception) and Group D (ART conception). We do not impose this strong assumption though. Instead, we make a conditional exogeneity/independence claim. This means assuming that the treatment assignments are independent of our outcomes of interest after controlling for the relevant confounders—variables causing changes in both treatment and outcome.
@@
We start from a descriptive analysis of our sample. Table 2 reports difference in means across treatment groups. We find a clinically important difference in maternal age, which is higher by 1.9 years (36.6-34.7 years) in mothers that spontaneously conceived after a failed ART —at an age where maternal fertility is in significant decline. This is also reflected in a lower maternal year of birth. In our sample, mothers who spontaneously conceived after ART are less likely to be at their first cycle ($0.25-0.36=-0.10$ p.p.) and have had more unsuccessful cycles before their pregnancy ($3.1-1.6=1.5$ cycles). A last noticeable difference is in the prevalence of male infertility, which is 7 p.p. higher in mothers of ART-conceived babies. The two groups are clinically similar in other respects, such as number of past pregnancies, socio-economic disadvantage, co-morbidities and cause of infertility.
Table 2: Covariate balance table
Control Treatment Difference
Variable Mean Obs. Mean Obs. Value p-value
Age 36.78 935 34.75 12169 2.024 0.0000
(5.13) (4.48) (0.163)
First cycle 0.26 935 0.36 12169 -0.099 0.0000
(0.44) (0.48) (0.015)
No. of past unsuccessful cycles 3.16 935 1.66 12169 1.501 0.0000
(2.53) (2.18) (0.073)
Maternal year of birth 1975.81 935 1977.98 12169 -2.166 0.0000
(5.21) (4.66) (0.167)
SEIFA 3.49 935 3.49 12169 0.001 0.9776
(1.46) (1.41) (0.048)
Remoteness: major city 0.9 935 0.89 12169 0.011 0.2843
(0.3) (0.31) (0.011)
Remoteness: inner regional 0.07 935 0.09 12169 -0.014 0.1058
(0.26) (0.28) (0.01)
Remoteness: outer regional 0.02 935 0.02 12169 0.005 0.3292
(0.15) (0.13) (0.005)
Remoteness: remote and very remote 0 935 0 12169 -0.001 0.2863
(0.03) (0.05) (0.001)
Mother born in Australia 0.61 935 0.66 12169 -0.053 0.0013
(0.49) (0.47) (0.016)
Had past pregnancies 0.4 935 0.32 12169 0.086 0.0000
(0.49) (0.47) (0.017)
No. of past pregnancies 0.53 935 0.37 12169 0.154 0.0000
(0.77) (0.62) (0.027)
Diabetes mellitus 0.01 935 0.01 12169 0.002 0.5584
(0.12) (0.11) (0.004)
Gestational diabetes 0.12 935 0.09 12169 0.033 0.0028
(0.33) (0.29) (0.011)
Chronic hypertension 0.01 935 0.01 12169 0.002 0.5394
(0.11) (0.1) (0.003)
Pre-eclampsia 0.02 935 0.02 12169 -0.001 0.8918
(0.14) (0.14) (0.004)
Gestational hypertension 0.04 935 0.05 12169 -0.008 0.2416
(0.2) (0.22) (0.007)
Smoked during pregnancy 0.02 935 0.01 12169 0.008 0.0996
(0.14) (0.11) (0.004)
Endometriosis 0.11 935 0.11 12169 -0.008 0.4604
(0.31) (0.32) (0.011)
Other cause of infertility 0.38 935 0.37 12169 0.011 0.4964
(0.49) (0.48) (0.016)
Male infertility 0.35 935 0.43 12169 -0.079 0.0000
(0.48) (0.49) (0.016)
Unexplained infertility 0.34 935 0.29 12169 0.055 0.0006
(0.48) (0.45) (0.016)
  • Notes: This table compares the covariate means of our treatment group—all ART births—and control group—spontaneous births conceived 3-to-12 months after a failed ART cycle. Each mean’s standard errors is reported below it between parentheses. Taking the difference between the reported control and treated means typically returns a value inconsistent with the reported difference. This is due to rounding.
@@ Results
Table 3 reports the OLS estimates of Equation 1 under our full sample (column 1), and under further sample restrictions (columns 2-6)  [^11] When restricting the sample, we adjust the model accordingly, so that, for instance, we drop the Number of previous pregnancies variable if we are conditioning on having no previous pregnancies.. In column 2, we only include births whose mother had no previous pregnancies longer than 20 weeks of gestation (nulliparous) nor previous unsuccessful ART cycles. In column 3, we only include births whose mother had no previous pregnancies longer than 20 weeks of gestation. In column 4, we only include, in the control group, births whose mother had a single previous unsuccessful ART cycles, and in the treatment group, ART births conceived at the first ART cycle. Finally, in column 5 we drop all our covariates and instead stratify by our most important covariates. More precisely, we estimate a difference in means across treatment groups in a sample restricted to nulliparous mothers at their first ART cycle, aged less than 40 and without comorbidities (selected primips).
Table 3: OLS: Effect of IVF conception on perinatal outcomes
(1) (2) (3) (4) (5) (6)
Spontaneous labour 0.042 -0.017 0.028 0.008 0.028 -0.017
(0.016) (0.041) (0.02) (0.031) (0.025) (0.054)
[0.006] [0.668] [0.157] [0.786] [0.263] [0.756]
Preterm spont. labour 0.009 0.01 0.014 0.006 0.014 0.006
(0.006) (0.014) (0.008) (0.01) (0.01) (0.019)
[0.156] [0.481] [0.075] [0.587] [0.157] [0.751]
C-section -0.035 0.046 -0.016 0.014 0.001 -0.006
(0.017) (0.04) (0.021) (0.031) (0.026) (0.053)
[0.039] [0.252] [0.437] [0.657] [0.964] [0.915]
Gestational age -0.077 -0.154 -0.145 -0.07 -0.222 -0.086
(0.086) (0.193) (0.12) (0.151) (0.14) (0.285)
[0.375] [0.423] [0.225] [0.642] [0.113] [0.764]
Pre-term birth 0.01 0.027 0.017 0.017 0.022 0.015
(0.01) (0.02) (0.013) (0.016) (0.015) (0.027)
[0.286] [0.179] [0.192] [0.294] [0.152] [0.58]
APGAR1 -0.001 0.202 0.02 0.134 0.002 0.364
(0.057) (0.154) (0.078) (0.109) (0.091) (0.211)
[0.983] [0.192] [0.795] [0.218] [0.981] [0.085]
APGAR5 0.012 0.149 -0.017 0.126 -0.031 0.306
(0.041) (0.108) (0.055) (0.081) (0.065) (0.168)
[0.764] [0.169] [0.756] [0.117] [0.634] [0.068]
Birth weight 24.057 -25.465 10.402 13.195 -8.251 15.619
(20.749) (47.036) (27.358) (36.889) (32.334) (67.219)
[0.246] [0.588] [0.704] [0.721] [0.799] [0.816]
Obs. by treatment 935/12169 153/3493 558/8311 268/4717 376/7199 91/2661
  • Notes: This table reports average treatment effect estimates for the effect of an IVF conception on perinatal outcomes, estimated via OLS. Each column reports estimates for a different model specification. Each row reports effects for a different outcome. Column (1) reports the preferred estimates, which use all sample and confounders. Column (2): baseline if preterm equal 1. Column (3): first pregnancy and cycle. Column (4): first pregnancy. Column (5) first cycle. Column (6) restricts the sample to women at their first pregnancy and of less than 40 years of age. Standard errors are reported between parentheses, while p-values between square brackets.
@@
The estimated treatment effects are, in general, clinically small. Focusing on our baseline estimates (Column 1), we find that ART has no significant effect on the rate of preterm births, on the rate of spontaneous preterm births, and on the average gestational age at birth—neither at the statistical nor at the clinical level. Taking preterm birth as an example, the 0.008 coefficient, with a 0.01 standard error, indicates that our best estimate for the effect of ART treatment of preterm birth is 0.8 p.p., but we cannot exclude that the true estimate is equal to zero at the 95% confidence level. The implied confidence interval is [$-0.0116,0.0276$], so that we cannot exclude effects ranging from -1.16 p.p. to +2.76 p.p [^12] Confidence intervals, while omitted in Table 3 due to space constraints, are reported for the double machine learning estimates, which are very similar to the OLS estimates and are reported Tables 4 and 5.
@@
The same applies to birth weight and to the APGAR 1 and 5 scores, whose estimated effects are not statistically significant and are clinically small. To give an idea of the magnitudes of these statistically insignificant estimates, the effect of ART on APGAR1 is estimated to be $-0.013$ points, on a 1-to-5 scale, while the effect on gestational age at birth is estimated to be $-0.057$ weeks, which less than 10 hours. As shown by the p-values in the square brackets, these estimates are not statistically significant at the 1% level or below, despite small standard errors. Considering that gestational age is measured in weeks, the APGAR scores are integers between zero and five and that birth weight is measured grams, we can notice how imposing additional restrictions on the sample does not shift the estimates in a clinically significant way for the most part. Understandably, our estimates’ standard errors and statistical significance varies as the sample gets smaller.
@@
Table 4 reports the baseline DML treatment effect estimates, while Table 5 reports the DML estimates after trimming the propensity score at 0.15 and 0.85. Not only the baseline DML estimates are robust to trimming, they are also virtually identical to the OLS estimates (Table 3 , Column 1), as can be visualised in Figures 7 and 6.
Table 4: DML: ATE estimates of ART conception on obstetric outcomes
Coef. s.e. l-CI r-CI p-value Obs.
Spontaneous labour 0.035 0.012 0.011 0.058 0.0034 13101
Preterm spont. labour 0.005 0.005 -0.004 0.014 0.2453 13101
C-section -0.028 0.013 -0.053 -0.003 0.0258 13104
Gestational age -0.038 0.062 -0.160 0.084 0.5413 13104
Pre-term birth 0.007 0.007 -0.008 0.021 0.3608 13104
APGAR1 0.028 0.041 -0.051 0.108 0.4865 13085
APGAR5 0.035 0.030 -0.023 0.093 0.2404 13087
Birth weight 18.953 15.248 -10.932 48.838 0.2139 13089
  • Notes: This tables reports our double machine learning estimates for the average treatment effect (ATE) of ART on obstetric outcomes. For each outcome (Column 1), we report the point estimate of the ATE (Column 2), its standard error (Column 3), its left and right 95% confidence interval extrema (Columns 4 and 5), its p-value (Column 6) and the number of observations (Column 7).
Table 5: DML: ATE estimates of ART conception on obstetric outcomes with trimming at 0.15
Coef. s.e. l-CI r-CI p-value Obs.
Spontaneous labour 0.034 0.012 0.010 0.057 0.0050 13101
Preterm spont. labour 0.005 0.005 -0.004 0.014 0.2983 13101
C-section -0.026 0.013 -0.051 -0.001 0.0418 13104
Gestational age -0.030 0.063 -0.153 0.093 0.6294 13104
Pre-term birth 0.007 0.007 -0.008 0.021 0.3622 13104
APGAR1 0.036 0.041 -0.046 0.117 0.3899 13085
APGAR5 0.035 0.030 -0.024 0.093 0.2488 13087
Birth weight 18.621 15.402 -11.567 48.809 0.2267 13089
  • Notes: This tables reports our double machine learning estimates for the average treatment effect (ATE) of ART on obstetric outcomes, after trimming births with propensity scores under 0.15 and above 0.85. For each outcome (Column 1), we report the point estimate of the ATE (Column 2), its standard error (Column 3), its left and right 95% confidence interval extrema (Columns 4 and 5), its p-value (Column 6) and the number of observations (Column 7).